If there are 2.38 x 1023 hydrogen atoms available, how many grams of dinitrogen tetrahydride could be made?
1
Expert's answer
2019-10-17T07:31:57-0400

Molecular weight of N2H4 is = 32 gm/mole

In 1 mole of N2H4 there are 4*6.023*1023 atoms of H is present.

2.38*1023 hydrogen atoms =(2.38*1023)/(4*6.023*1023) moles of N2H4

=101.6*1023 moles of N2H4

Therefore 101.6*1023 moles = 32*101.6*1023 gm of N2H4 can be formed